answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Who started the fidget spinner company?

"Catherine Hettinger has an inventor's mind: When she notices a problem, she tries to imagine a toy or device that can remedy it.

So when she heard about young boys throwing rocks at police officers and people walking past them while visiting her sister in Israel, her wheels started turning. She started brainstorming devices that could distract young children and provide them with a soothing toy to play with.

First, she thought of a soft rock that kids could throw. But then she tossed that idea aside, still thinking about other options when she returned to her home in Orlando, Fla. It was there that she eventually developed the idea that would become the original fidget spinner - more than two decades before the wildly popular device became the must-have toy for both kids and adults this year."

-h'ttp://time.com/money/4762207/fidget-spinner-inventor-catherine-hettinger

What is the probability of choosing a red marble or a green marble out of a bag containing 5 blue marbles 3 red marbles and 1 green marble?

Pr = ways_of_success/total_ways

Pr(Red) = 3/(5+3+1) = 3/9

Pr(Green) = 1/9

As they are independent:

Pr(Red or Green) = Pr(Red) + Pr(Green) = 3/9 + 1/9 = 4/9

----------------------------------------------------

Alternatively, as there are only three colours, selecting either Red or Green is the same as not selecting Blue:

Pr(Red or Green) = Pr(not Blue) = 1 - Pr(Blue)

→ Pr(Red or Green) = 1 - 5/9 = 4/9

In each case are the two events dependent or independent Toss a dime and then toss a quarter. Pick a card from a deck of cards. Keep the card and pick another.?

Picking cards without replacing them are NOT independent events, since once you pick a card, you can't pick the same card again.

Are mutually exclusive events?

Mutually exclusive events are occurrences where, say, a couple of propositions are possible, but if one occurs, the other cannot. A coin toss might be a good example. A coin lands heads or it lands tails. It cannot land on both in the same toss. A coin toss, therefore, can be said to be a mutually exclusive event.

What is true of a random sample?

Every member in the population has the same probability of being in the sample.Or, equivalently, every set of the given sample size has the same probability of being selected.

What is the experimental probability that the next toss and spin will result in 3 and tails?

The experimental probability, by definition, can only be determined after you have carried out the experiment!

What is the probability of pulling a king or queen?

The answer depends on whether or not the card is being pulled at random and also form what set of cards.

What is the probability of choosing a vowel in the alphabet?

The answer depends on

  • the alphabet that you chose,
  • whether or not you consider y to be a vowel,
  • whether or not each letter is equally likely to be chosen (eg not from a bag of scrabble tiles), and
  • whether or not the choice was random.

If a choice was random, from equal numbers of letters from the modern Roman alphabet and y is not considered a vowel then the answer is 5/26.

The probability of obtaining a sum greater 8 or less than 6?

The answer depends on what the experiment consists of. Since that is not specified, there can be no sensible answer.

Which of these is best used for displaying frequency distributions when one or two categories make up a large proportion of the total?

The answer depends on what is meant by two or more. If it refers to around 4 to 7 then a pie chart may be best: it depends on whether it is appropriate to lump all the remaining categories as an "other" category.

If there are 6 or more categories then a bar chart may be appropriate but again, that depends on whether lumping the remainder together as an "other" category makes sense.